OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (5, –9); m = -2

OpenStudy (amistre64):

point slope eh...

OpenStudy (anonymous):

(y+9)/(x-5)=-2 y+9= -2x+10

OpenStudy (amistre64):

y = m(x-Px)+Py y-Py = m(x-Px) is the point slope form

OpenStudy (amistre64):

y+9 = -2(x-5)